Фn undergraduate student weighs 20g of sodium carbonate salt. Of Na =23, O = 16, Carbon = 12 and H = 1, what is the mole of the hydroxide?
1
Expert's answer
2015-11-13T11:33:14-0500
Solution
M(NaOH) = 23 + 16 + 1 = 40 (g/mol)
n(NaOH) = m/M = 20/40 = 0.5 (mol)
Answer
n(NaOH) = 0.5 mol